hallm:

15. You want to calculate the displacement of an object thrown over a bridge. Using -10 m/s2 for acceleration due to gravity, what would be the total displacement of the object if it took 8 seconds before hitting the water?

mhchen:

Do you know the formula \[x_{final} = x_{initial}+v_{initial}(t) + \frac{1}{2}a^2\] The initial position is at 0. The initial velocity is also at 0. It says to use -10m/s^2 for acceleration. It also says it took 8 seconds to hit the water. Plug that into the equation we get: \[x = 0+0(8) + \frac{1}{2}(-10)^2\] Can you solve for x here? Once you solve that displacement has a direction, and since the direction the ball is falling is downwards, it would be negative.

sillybilly123:

why is the initial velocity = 0 ? "dropped", maybe works. and the sign convention - is upwards or downwards positive i think those might be the more important questions

mhchen:

Because before you throw it the ball isn't moving so initial velocity is 0 Upwards is positive downwards is negative

sillybilly123:

yes Qu: why is downwards negative?

mhchen:

Basically in an XY coordinate plane going downwards is negative.

sillybilly123:

for every quantity ?

mhchen:

If you're talking about an object moving then yes

sillybilly123:

can't we define acceleration as upward?

mhchen:

If the ball was thrown upwards then yeah it'll be going positive. But gravity pulls the ball downwards

sillybilly123:

forgive me: but on earth, you can only ever have one co-ordinate system for kinematic problems?

sillybilly123:

i think that you're really talking about vectors